### v2

# How to explain grokking

Abstract
Explanation of grokking (delayed generalization) in learning is given by modeling grokking by the stochastic gradient Langevin dynamics (Brownian motion) and applying the ideas of thermodynamics.
1 Learning and thermodynamics
1.1 SGLD and Eyring’s formula
Gradient descent is the numerical solution of the differential equation
$\frac{dx}{dt}=-f^{\prime}(x),$
the trajectory goes to the local minimum of $f$.
For the numerical iteration of the descent, the vector $x$ will change as
$x_{k+1}=x_{k}-\alpha_{k}f^{\prime}(x_{k}).$
The stochastic gradient Langevin dynamics (or SGLD) is given by iteration of
$x_{k+1}=x_{k}+w_{k}-\alpha_{k}f^{\prime}(x_{k}),$
where $w_{k}$ are independent Gaussian random vectors.
Continuous limit of SGLD is described by the stochastic differential equation (SDE) for Brownian motion in a potential
$d\xi^{i}(t)=\sqrt{2\theta}dw^{i}(t)-{\partial f(\xi(t))\over\partial x^{i}}dt,$ (1)
where $dw^{i}(t)$ is the stochastic differential of Wiener process which satisfies
$dw^{i}(t)dw^{j}(t)=\delta_{ij}dt.$
This equation was discussed in relation to learning in [1], [2].
The Fokker–Planck equation for SGLD (1) reduces to the diffusion equation in the potential
${\partial u\over\partial t}=\theta\Delta u+\nabla u\cdot\nabla f+u\Delta f,$ (2)
where $x\in\mathbb{R}^{d}$, $u=u(x,t)$ is the distribution function, $f=f(x)$ is the potential, $f\in C^{2}(\mathbb{R}^{d})$, $\theta>0$ is the temperature. Equivalently,
${\partial u\over\partial t}=\theta\,{\rm\bf div}\,\left[e^{-\beta f}\,{\rm\bf grad}\,\left[ue^{\beta f}\right]\right],$
therefore the Gibbs distribution $e^{-\beta f}$, $\beta=1/\theta$ is a stationary solution of the equation, and the solution converges to the Gibbs distribution (under some conditions on $f$).
The Eyring formula of kinetic theory describes the reaction rate (transition between two potential wells due to diffusion of the type (2)): the reaction rate is proportional to
$e^{-\beta(F_{1}-F_{0})},$ (3)
where $F_{1}$ is the free energy of the transition state (the saddle between two potential wells) and $F_{0}$ is the free energy of the initial state of the reaction (the potential well from which the transition occurs).
The free energy of a state is $F=E-\theta S$, where $E$ is the energy and $S$ is the entropy of the state, in the general case
[…]
High entropy of the potential well lowers the free energy, that is, wide potential wells better capture the particle.
1.2 SGLD and learning
Let a training sample $\{z_{l}\}$, $l=1,\dots,L$ and a loss function $\mathcal{L}(z,x)\geq 0$ for test $z$ and hypothesis $x$ (let the hypothesis space be $\mathbb{R}^{d}$) be given.
Minimizing the empirical risk (learning) is the problem
$f(\{z\},x)=\frac{1}{L}\sum\limits_{l=1}^{L}\mathcal{L}(z_{l},x)\,\to\,\min\limits_{x}.$ (4)
We will assume $\mathcal{L}(z,x)\geq 0$, $\forall z$, hence $f(\{z\},x)\geq 0$.
Overfitting is the failure to generalize to a learning problem (4) when the sample $\{z\}$ is replaced (in particular, low risk for training sample, high risk for validation sample).
We will use the known approach of flat minima: narrow (sharp) minima of empirical risk (in the hypothesis space) are associated with overfitting, and wide (flat) minima correspond to solutions of the learning problem with generalization [3]. This known observation is related to algorithmic stability, i.e. stability of the solution of a learning problem to perturbations of the training sample [4], [5], [6].
For the learning problem (4) we will use the SGLD algorithm (1) in the hypothesis space, equivalently, the diffusion equation in the potential (2). Therefore, the distribution function $u(x)$ converges to the Gibbs distribution concentrated in potential wells with low free energy.
Then Eyring’s (3) reaction rate formula predicts the capture of the SGLD learning result by wide potential wells (due to the entropic part of free energy), i.e., the reduction of the overfitting effect in the flat minima approach. This explains the effect of reducing of overfitting for SGLD. We believe the same effect will work for other forms of stochastic gradient descent (such as mini-batch procedure). This (and other) approaches to control overfitting are discussed in [7].
2 Grokking and thermodynamics
Grokking (delayed generalization) phenomenon in learning was discovered in [8], see also [9], [10], [11]. Overparameterized network was trained to perform modular arithmetics (in particular, addition and other operations modulo 97). The phenomenon of delayed generalization was found, and the following observations were made:
1) The model was trained in $10^{3}$ steps (by the stochastic gradient descent), memorization of training sample was achieved with almost $100\%$ error on the validation set (i.e. total overfitting). If the training was continued, then in about $10^{6}$ steps total generalization was obtained (no overfitting).
2) Exponential growth of grokking time with decreasing of training sample was observed.
3) Some kind of principal component analysis for vector representations (embeddings) of residues in the obtained model of modular arithmetics was performed. It was observed that embeddings of residues lie approximately on a circle, adding a residue is a shift along such a circle. Therefore, modular addition, performed by the neural network, looks like ”structure” (an algorithm, given by trigonometric polynomials [10]).
We propose the following explanation of grokking.
For overparameterized models local minima merge into a manifold of zero empirical risk in the hypothesis space [12], given by the condition (since the loss function is non-negative)
$f(\{z\},x)=0,\quad{\rm i.e.}\quad\mathcal{L}(z,x)=0,\,\forall z\in\{z\}.$
The zero-risk manifold contains narrows (with low entropy) and wide valleys (with high entropy). The correct solution for learning problem (”structure”, for example algorithm of modular addition) mostly likely lies in the high-entropy region of the zero-risk manifold.
Memorization of the training sample is achieving the zero risk manifold $f(\{z\},x)=0$ by the stochastic gradient descent with non-zero gradient, in this case the path traveled in the hypothesis space is proportional to the number of descent steps tsimilar-$\sim t$.
Grokking is a random walk in the zero risk manifold $f(\{z\},x)=0$ (or Brownian motion). For Brownian motion the path traveled is proportional to the root of the number of steps tsimilar-$\sim\sqrt{t}$. By the second law of thermodynamics the system will travel to areas of high entropy at the zero risk manifold.
If we assume that the memorization and grokking result in similar path lengths in the zero risk manifold, then the duration of grokking (in the SGD steps) will be the square of the duration of memorization, which is observed in simulation ($10^{3}$ and $10^{6}$).
As the training sample $\{z\}$ grows, the zero-risk manifold $f(\{z\},x)=0$ shrinks since additional conditions $\mathcal{L}(z,x)=0$ for $z\in\{z\}$ are imposed. A solution of the learning problem in the form of an algorithm (”structure”) exists for any training sample; therefore, as the training sample grows, the zero-risk manifold shrinks toward a region with high entropy containing the desired solution in the form of an algorithm.
Let us suppose that the imposition of each additional condition $\mathcal{L}(z,x)=0$, $z\in\{z\}$ removes an equal percentage of the volume of the zero-risk manifold as the training sample size increases. Then the entropy of the zero-risk region decreases linearly as the sample size increases.
Example: the zero-risk manifold is (initially) a multidimensional cube $[0,L]^{d}$ with side $L$. Imposing the condition $\mathcal{L}(z,x)=0$, $z\in\{z\}$ is imposing the constraint $0\leq x\leq\varepsilon L$, $\varepsilon<1$ on one of the coordinates $x$ of the cube. Then, choosing $z\in\{z\}$ at each step and imposing the constraint on the corresponding coordinate, we obtain an exponential decrease in the volume of the zero-risk manifold with increasing sample size (and linear decrease in entropy).
Grokking is the transition from the initial region (part of the zero-risk manifold where the Brownian motion of grokking starts) to the grokking region (a valley with high entropy, a neighborhood of the solution of the learning problem in the form of a ”structure”).
The Eyring’s formula $e^{-\beta(F_{1}-F_{0})}$ gives an approximation of the dependence of the reciprocal of the grokking time on the training sample size. Here $F_{0}$ and $F_{1}$ are free energies of the initial region for grokking and of the transition state (the neighborhood of the solution). The entropy of the initial region decreases linearly with the training sample size (hence the free energy $F_{0}=E_{0}-\theta S_{0}$ increases linearly). The free energy $F_{1}$ of the transition region changes slightly with the training sample size. Consequently, the grokking time increases exponentially when the training sample decreases (this agrees with observations in [8]).
3 Summary
Simple ideas of thermodynamics and kinetic theory allow us to explain better generalization observed for learning by the stochastic gradient optimization procedure, see also [7] (where also overfitting control for GAN model was discussed).
We also have explained the grokking (delayed generalization) phenomenon and some properties of grokking observed in [8].