### v3

# How to explain grokking

Abstract
Explanation of grokking (delayed generalization) in learning is given by modeling grokking by the stochastic gradient Langevin dynamics (Brownian motion) and applying the ideas of thermodynamics.
1 Introduction
Stochastic gradient descent, which is a basic model in machine learning, when taken in the form of the stochastic gradient Langevin dynamics, is a direct analogue of a random walk along the reaction coordinates, which is a principal model in chemical kinetics. It is natural to use in learning the ideas of kinetics and thermodynamics, in particular, if the gradient descent is an analogue of the energy driven chemical reaction, then the stochastic gradient descent should include also the discussion of entropy, i.e. of the free energy minimization. We show that the transition to generalization in grokking is an analogue of the entropy driven chemical reaction (see the discussion below).
2 Learning and thermodynamics
2.1 SGLD and Eyring’s formula
Gradient descent is the numerical solution of the differential equation
$\frac{dx}{dt}=-f^{\prime}(x),$
the trajectory goes to the local minimum of $f$.
For the numerical iteration of the descent, the vector $x$ will change as
$x_{k+1}=x_{k}-\alpha_{k}f^{\prime}(x_{k}).$
The stochastic gradient Langevin dynamics (or SGLD) is given by iteration of
$x_{k+1}=x_{k}+w_{k}-\alpha_{k}f^{\prime}(x_{k}),$
where $w_{k}$ are independent Gaussian random vectors.
Continuous limit of SGLD is described by the stochastic differential equation (SDE) for Brownian motion in a potential
$d\xi^{i}(t)=\sqrt{2\theta}dw^{i}(t)-{\partial f(\xi(t))\over\partial x^{i}}dt,$ (1)
where $dw^{i}(t)$ is the stochastic differential of Wiener process which satisfies
$dw^{i}(t)dw^{j}(t)=\delta_{ij}dt.$
This equation was discussed in relation to learning in [1], [2].
The Fokker–Planck equation for SGLD (1) reduces to the diffusion equation in the potential
${\partial u\over\partial t}=\theta\Delta u+\nabla u\cdot\nabla f+u\Delta f,$ (2)
where $x\in\mathbb{R}^{d}$, $u=u(x,t)$ is the distribution function, $f=f(x)$ is the potential, $f\in C^{2}(\mathbb{R}^{d})$, $\theta>0$ is the temperature. Equivalently,
${\partial u\over\partial t}=\theta\,{\rm\bf div}\,\left[e^{-\beta f}\,{\rm\bf grad}\,\left[ue^{\beta f}\right]\right],$
therefore the Gibbs distribution $e^{-\beta f}$, $\beta=1/\theta$ is a stationary solution of the equation, and the solution converges to the Gibbs distribution (under some conditions on $f$).
Remark. Let us discuss a short introduction to statistical mechanics for mathematicians, based on arguments from [3]. Let us put (2) as
${\partial u\over\partial t}=-\theta e^{-\beta f/2}A^{*}Ae^{\beta f/2}u,\quad A=e^{-\beta f/2}\nabla e^{\beta f/2}.$
The operator $A^{*}A$ is positive, moreover
$e^{-t\theta e^{-\beta f/2}A^{*}Ae^{\beta f/2}}=e^{-\beta f/2}e^{-t\theta A^{*}A}e^{\beta f/2},$
$e^{-t\theta A^{*}A}$ is a contraction semigroup in $L_{2}$. The condition $e^{\beta f/2}u\in L_{2}$ is
[…]
For the stationary solution $e^{-\beta f}$ eq. (3) is the integrability of the Gibbs distribution
[…]
zero is the eigenvalue of $A^{*}A$
$\left[e^{-\beta f/2}A^{*}Ae^{\beta f/2}\right]e^{-\beta f}=0.$
Then if the initial condition $u_{0}$ satisfies (3) and Gibbs distribution is integrable (4), the dynamics $e^{-t\theta A^{*}A}e^{\beta f/2}u_{0}$ belongs to $L_{2}$ and converges to $e^{-\beta f/2}$, and the solution $e^{-\beta f/2}e^{-t\theta A^{*}A}e^{\beta f/2}u_{0}$ of (2) converges to Gibbs distribution $e^{-\beta f}$.
The Eyring formula of kinetic theory describes the reaction rate (for the transition between two potential wells due to diffusion of the type (2)): the reaction rate is proportional to
$e^{-\beta(F_{1}-F_{0})},$ (5)
where $F_{1}$ is the free energy of the transition state (the saddle between two potential wells) and $F_{0}$ is the free energy of the initial state of the reaction (the potential well from which the transition occurs).
The free energy of a state is $F=E-\theta S$, where $E$ is the energy and $S$ is the entropy of the state, in the general case
[…]
High entropy of the potential well lowers the free energy, that is, wide potential wells better capture the particle for the dynamics (2).
2.2 SGLD and learning
Let a training sample $\{z_{l}\}$, $l=1,\dots,L$ and a loss function $\mathcal{L}(z,x)\geq 0$ for test $z$ and hypothesis $x$ (let the hypothesis space be $\mathbb{R}^{d}$) be given.
Minimizing the empirical risk (learning) is the problem
$f(\{z\},x)=\frac{1}{L}\sum\limits_{l=1}^{L}\mathcal{L}(z_{l},x)\,\to\,\min\limits_{x}.$ (6)
We will assume $\mathcal{L}(z,x)\geq 0$, $\forall z$, hence $f(\{z\},x)\geq 0$.
Overfitting is the failure to generalize to a learning problem (6) when the sample $\{z\}$ is replaced (in particular, low risk for training sample, high risk for validation sample).
We will use the known approach of flat minima: narrow (sharp) minima of empirical risk (in the hypothesis space) are associated with overfitting, and wide (flat) minima correspond to solutions of the learning problem with generalization [4]. This known observation is related to algorithmic stability, i.e. stability of the solution of a learning problem to perturbations of the training sample [5], [6], [7].
For the learning problem (6) we will use the SGLD algorithm (1) in the hypothesis space, equivalently, the diffusion equation in the potential (2). Therefore, the distribution function $u(x)$ converges to the Gibbs distribution concentrated in potential wells with low free energy.
Then Eyring’s (5) reaction rate formula predicts the capture of the SGLD learning result by wide potential wells (due to the entropic part of free energy), i.e., the reduction of the overfitting effect in the flat minima approach. This explains the effect of reducing of overfitting for SGLD. We believe the same effect will work for other forms of stochastic gradient descent (such as mini-batch procedure). This (and other) approaches to control overfitting are discussed in [8].
Therefore, we consider the stochastic gradient optimization as a minimization of free energy of a potential well in statistical mechanics, or an analogue of chemical reaction.
Gradient descent optimization is a chemical reaction by energy minimization, but in general also entropic contribution to free energy is important.
3 Grokking and thermodynamics
Grokking (delayed generalization) phenomenon in learning was discovered in [9], see also [10], [11], [12], [13]. Overparameterized network was trained to perform modular arithmetics (in particular, addition and other operations modulo 97). The phenomenon of delayed generalization was found, and the following observations were made [9]:
1) The model was trained in $10^{3}$ steps (by the stochastic gradient descent), memorization of training sample was achieved with almost $100\%$ error on the validation set (i.e. total overfitting). If the training was continued, then in about $10^{6}$ steps total generalization was obtained (no overfitting).
2) Exponential growth of grokking time with decreasing of training sample was observed.
3) Some kind of principal component analysis for vector representations (embeddings) of residues in the obtained model of modular arithmetics was performed. It was observed that embeddings of residues lie approximately on a circle, adding a residue is a shift along such a circle. Therefore, modular addition, performed by the neural network, looks like ”structure” (an algorithm, given by trigonometric polynomials [12]).
We propose the following explanation of grokking.
For overparameterized models local minima merge into a manifold of zero empirical risk in the hypothesis space [14], given by the condition (since the loss function is non-negative)
$f(\{z\},x)=0,\quad{\rm i.e.}\quad\mathcal{L}(z,x)=0,\,\forall z\in\{z\}.$
The zero-risk manifold contains narrows, or ravines (areas with low entropy) and wide valleys (with high entropy). For ravine, or river–valleys landscape, learning is a motion along the ravine (the reaction coordinate in chemistry). The correct solution for learning problem (”structure”, for example algorithm of modular addition) mostly likely lies in the high-entropy region of the zero-risk manifold. The ravine method [15] was developed to improve optimization for such a landscape.
Memorization of the training sample is achieving the zero risk manifold $f(\{z\},x)=0$ by the stochastic gradient descent with non-zero gradient, in this case the path traveled in the hypothesis space is proportional to the number of descent steps tsimilar-$\sim t$.
Grokking is a random walk in the zero risk manifold $f(\{z\},x)=0$ (or Brownian motion). For Brownian motion the path traveled is proportional to the root of the number of steps tsimilar-$\sim\sqrt{t}$.
By the second law of thermodynamics the system will travel to areas of high entropy at the zero risk manifold, i.e. the transition to generalization in grokking is an analogue of the entropic mechanism of chemical reaction (moreover the energy difference for different points in the zero risk manifold equals to zero).
If we assume that the memorization and grokking result in similar path lengths in the zero risk manifold, then the duration of grokking (in the SGD steps) will be the square of the duration of memorization, which is observed in simulation ($10^{3}$ and $10^{6}$).
As the training sample $\{z\}$ grows, the zero-risk manifold $f(\{z\},x)=0$ shrinks since additional conditions $\mathcal{L}(z,x)=0$ for $z\in\{z\}$ are imposed. A solution of the learning problem in the form of an algorithm (”structure”) exists for any training sample; therefore, as the training sample grows, the zero-risk manifold shrinks toward a region with high entropy containing the desired solution in the form of an algorithm.
Let us suppose that the imposition of each additional condition $\mathcal{L}(z,x)=0$, $z\in\{z\}$ removes an equal percentage of the volume of the zero-risk manifold as the training sample size increases. Then the entropy of the zero-risk region decreases linearly as the sample size increases.
Example: the zero-risk manifold is (initially) a multidimensional cube $[0,L]^{d}$ with side $L$. Imposing the condition $\mathcal{L}(z,x)=0$, $z\in\{z\}$ is imposing the constraint $0\leq x\leq\varepsilon L$, $\varepsilon<1$ on one of the coordinates $x$ of the cube. Then, choosing $z\in\{z\}$ at each step and imposing the constraint on the corresponding coordinate, we obtain an exponential decrease in the volume of the zero-risk manifold with increasing sample size (and linear decrease in entropy).
Grokking is the transition from the initial region (part of the zero-risk manifold where the Brownian motion of grokking starts) to the grokking region (a valley with high entropy, a neighborhood of the solution of the learning problem in the form of a ”structure”) along the ”reaction coordinate” in the zero-risk manifold.
The Eyring’s formula $e^{-\beta(F_{1}-F_{0})}$ gives an approximation of the dependence of the reciprocal of the grokking time on the training sample size. Here $F_{0}$ and $F_{1}$ are free energies of the initial region for grokking and of the transition state (the neighborhood of the solution). The entropy of the initial region decreases linearly with the training sample size (hence the free energy $F_{0}=E_{0}-\theta S_{0}$ increases linearly). The free energy $F_{1}$ of the transition region changes slightly with the training sample size. Consequently, the grokking time decreases exponentially when the training sample increases (this agrees with observations in [9]).
Remark: Transition State Theory in catalysis and Grokking
Catalysis in chemistry works by increasing entropy $S_{1}$ and hence lowering free energy $F_{1}=E_{1}-\theta S_{1}$ of the transition state in Eyring’s formula $e^{-\beta(F_{1}-F_{0})}$ of the reaction rate.
For grokking — when the training sample increases, entropy of the initial state $S_{0}$ decreases and free energy $F_{0}=E_{0}-\theta S_{0}$ increases.
Both catalysis and grokking depend on the free energy difference $F_{1}-F_{0}$.
Therefore, increasing of the sample size in grokking works as catalysis in entropy driven chemical reaction.
4 Summary
Simple ideas of thermodynamics and kinetic theory allow us to explain better generalization observed for learning by the stochastic gradient optimization procedure, see also [8] (where also overfitting control for GAN model was discussed).
We also have explained the grokking (delayed generalization) phenomenon and some properties of grokking observed in [9] as a manifestation of thermodynamic and kinetic phenomena — transition to generalization in grokking proceeds by entropic mechanism of chemical reaction, properties of grokking are explained by Brownian motion and Eyring’s formula. Delay of generalization in grokking is a result of transition through the barrier of entropy, and the delay is large when entropy of the initial state and therefore the barrier of free energy in Eyring’s formula is large.